# From Complex-Analytic Models to Sparse Domination: A Dyadic Approach of Hypersingular Operators via Bourgain’s Interpolation Method

### v3

@@ 4. Critical line estimates for the hypersingular Bergman projection
which can be regarded as a singular integral counterpart of the hypersingular maximal operator $M_{t}^{{\mathcal{D}}}$. We first recall that the off-critical line $L^{p}$ theory was studied in [12, Theorem 3]. In particular, they showed that for $K_{2t}$ is bounded from $L^{p}({\mathbb{D}})$ to $L^{q}({\mathbb{D}})$ if and only 1/q1/p>2t21/q-1/p>2t-2 for $1\leq p,q\leq\infty$, which is exactly the off-critical line regime for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$. Therefore, it is natural to consider the behavior of $K_{2t}$ on the critical line
@@
where in the last estimate, we used the standard integral estimate [31, Theorem 1.12]. Therefore, for any $\alpha>0$, as in Lemma 3.4,
@@
On the other hand, by [31, Theorem 1.12], one has
@@
First, recall that for any $z,w\in{\mathbb{D}}$, there exists a Carleson tent $Q_{J}$, containing both $z$ and $w$ such that $|Q_{J}|\simeq\left|1-z\overline{w}\right|^{2}$ (see, e.g., [25, Section 3]). Then using the $1/3$-trick, one can find some $I\in{\mathcal{D}}$ or $\widetilde{{\mathcal{D}}}$ (see, Remark 1.4), such that
@@
where ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ is the hypersingular sparse operator associated to a dyadic system $\mathcal{D}$} on $\mathbb{T}$, defined as [^4] We call ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ a hypersingular sparse operator since, on the critical line $\{(\tfrac{1}{p},\tfrac{1}{q}):\ \tfrac{1}{q}-\tfrac{1}{p}=2t-2\}$, it fails to satisfy strong $(p,q)$ bounds. Indeed, any strong-type estimate for ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ at a point on this line would imply the corresponding bound for $K_{2t}$, which is known to be false by [12, Theorem 3].
@@ 5.1. Graded family
    Indeed, $K_{{\mathcal{G}}}^{\prime}$ only controls the ratios $\mathfrak{G}_{j}/\mathfrak{G}_{j+1}$ for sufficiently large $j$ and, in particular, imposes no restriction on the initial scales (for instance, on $\mathfrak{G}_{0}/\mathfrak{G}_{1}$). Consequently, one loses uniform control on the gaps between the first few layers. Using the same idea as in Observation 1.8, one can construct a sequence of sparse families with the same sparseness and the same value of $K_{{\mathcal{G}}}^{\prime}$, but for which the associated hypersingular sparse operators still exhibit the “blow-up” phenomenon described in Example 1.8.
@@
Theorem 5.7 strengthens our earlier results on the boundedness behavior of $K_{2t}$ (see, Section 4). Moreover, the proof of Step II in Theorem 5.7 yields an alternative proof of [12, Theorem 3]. Finally, we note that the endpoint weak-type estimate (5.3) fails in general (see, Proposition 4.2).
@@ 6. Some open problems
where ${\mathbb{B}}_{n}$ is the unit ball in $\mathbb{C}^{n}$ and $dV$ is the normalized volume measure on ${\mathbb{B}}_{n}$. In [28], Zhao and Zhou characterized the strong-type $L^{p}_{\alpha}({\mathbb{B}}_{n})\to L^{q}_{\beta}({\mathbb{B}}_{n})$ bounds for these operators under various assumptions on the parameters $a,b,c,\alpha$, and $\beta$. Here, for $1\leq p<\infty$ and $1<\alpha<\infty$, the space $L^{p}_{\alpha}({\mathbb{B}}_{n})$ is defined with respect to the measure $dv_{\alpha}(z)=c_{\alpha}(1-|z|^{2})^{\alpha}\,dV(z)$, where $c_{\alpha}$ is chosen so that $v_{\alpha}({\mathbb{B}}_{n})=1$. A natural question is whether the method developed in the present paper can be applied to obtain critical-line estimates for Forelli–Rudin type operators in the hypersingular regime (see, e.g., [28, Theorem 1.1]).

### v4

@@ 2. $\bullet$
3.3. A two-weight estimate for $\mathcal{M}_{t}^{\mathcal{D}}$
The last part of this section concerns weighted estimates for $\mathcal{M}_{t}^{\mathcal{D}},\;t>0$ in the (hyper-singular) regime $\{1\leq p,q\leq+\infty:p>q\}$. We start with some definition.
Definition 3.9.
Let $\eta$ be a weight on $\mathbb{D}$} and $\mathcal{D}$} be a dyadic system on $\mathbb{T}$. We say $\eta$ belongs to the dyadic $\bf B$$_{\infty}$ class ${\bf B}_{\infty}({\mathcal{D}})$ if there exists an absolute constant $C>1$, such that for each $I\in{\mathcal{D}}$,
$\eta(Q_{I})\leq C\eta(Q_{I}^{\textrm{up}}).$
Remark 3.10.
The ${\bf B}_{\infty}(\mathcal{D})$ condition is a rather mild assumption, and it includes most of the well–known examples of weights on $\mathbb{D}$. Here are some examples.
  1. (1)
    Radial weights $\nu(z)=(1-|z|^{2})^{\alpha}$, $\alpha>-1$.
  2. (2)
    Békollé–Bonami ${\bf B}_{p}$ weights for $1<p<\infty$ (see, e.g., [10]).
  3. (3)
    Weights $\eta$ that satisfy both bounded hyperbolic oscillation, i.e., there exists $C_{\eta}>0$ such that for every arc $I\subseteq\mathbb{T}$,
    $C_{\eta}^{-1}\eta(\xi)\leq\eta(z)\leq C_{\eta}\,\eta(\xi),\qquad z,\xi\in Q_{I}^{\mathrm{up}},$
    and the Fuji–Wilson property
    $\sup\left\{\frac{\int_{Q_{I}}M(\eta\,\mbox{1\hskip-4.25pt{1}}_{Q_{I}})}{\int_{Q_{I}}\eta}:I\subseteq\mathbb{T}\right\}<+\infty$
    (see [1, Theorem 1.7]).
To this end, we refer the interested reader to the recent excellent papers [1, 28] and the references therein for a more systematic study of arbitrary Békollé–Bonami weights.
We have the following result.
Theorem 3.11.
Let $1\leq q<p\leq+\infty$ and $t>0$. Let $\mathcal{D}$ be a dyadic system on $\mathbb{T}$. Let $\mu$ and $\omega$ be two weights on $\mathbb{D}$, and set $\sigma:=\omega^{-1/(p-1)}$. Assume that $\mu,\sigma\in{\mathcal{B}}_{\infty}({\mathcal{D}})$. Then the following statements are equivalent.
  1. (1)
    ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ extends a bounded operator from $L^{p}(\omega,{\mathbb{D}})$ to $L^{q}(\mu,{\mathbb{D}})$;
  2. (2)
    Define
    $\phi(z):=\sum_{I\in{\mathcal{D}}}\beta_{I}\mbox{1\hskip-4.25pt{1}}_{Q_{I}^{\textrm{up}}}(z)$
    where
    $\beta_{I}:=\frac{1}{|Q_{I}|^{(t-1)q}}\cdot\frac{\mu(Q_{I})}{|Q_{I}|}\cdot\left(\frac{\sigma(Q_{I})}{|Q_{I}|}\right)^{\frac{q}{p^{\prime}}}.$
    Then $\phi\in L^{\frac{p}{p-q}}({\mathbb{D}})$.
Proof.
$(2)\Longrightarrow(1)$.   Assume (2). For any $\alpha>0$, as usual, let $E_{\alpha}:=\{z\in{\mathbb{D}}:{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)>\alpha\}$. Again, write $E_{\alpha}$ into disjoint union of maximal Carleson boxes $\{Q_{\alpha,i}\}_{i\geq 1}$ with
$\frac{1}{|Q_{\alpha,i}|^{t}}\int_{Q_{\alpha,i}}|f(z)|dA(z)>\alpha,\qquad\textrm{for}\quad i\geq 1.$
Then, using the assumption that $\sigma\in{\bf B}_{\infty}({\mathcal{D}})$,
$\alpha^{q}\mu(E_{\alpha})$ $=\alpha^{q}\sum_{i=1}^{\infty}\mu(Q_{\alpha,i})\lesssim\sum_{i=1}^{\infty}\frac{\mu(Q_{\alpha,i})}{|Q_{\alpha,i}|^{tq}}\left(\int_{Q_{\alpha,i}}|f(z)|dA(z)\right)^{q}$
$=\sum_{i=1}^{\infty}\frac{\mu(Q_{\alpha,i})}{|Q_{\alpha,i}|^{tq}}\left[\sigma(Q_{\alpha,i})\cdot\frac{1}{\sigma(Q_{\alpha,i})}\int_{Q_{\alpha,i}}|f(z)|\sigma^{-1}(z)\sigma(z)dA(z)\right]^{q}$
$\lesssim\sum_{i=1}^{\infty}\frac{\mu(Q_{\alpha,i})}{|Q_{\alpha,i}|^{tq}}\left[\sigma(Q^{\textrm{up}}_{\alpha,i})\cdot\frac{1}{\sigma(Q_{\alpha,i})}\int_{Q_{\alpha,i}}|f(z)|\sigma^{-1}(z)\sigma(z)dA(z)\right]^{q}$
(3.20) $\lesssim\sum_{i=1}^{\infty}\frac{\mu(Q_{\alpha,i})}{|Q_{\alpha,i}|^{tq}}\left[\int_{Q_{\alpha,i}^{\textrm{up}}}{\mathcal{M}}^{{\mathcal{D}},\sigma}\left(|f|\sigma^{-1}\right)(z)\sigma(z)dA(z)\right]^{q},$
where ${\mathcal{M}}^{{\mathcal{D}},\sigma}$ is the weighted dyadic maximal operator given by
${\mathcal{M}}^{{\mathcal{D}},\sigma}f(z):=\sup_{z\in Q_{I},\;I\in{\mathcal{D}}}\frac{1}{\sigma(Q_{I})}\int_{Q_{I}}|f(z)|\sigma(z)dA(z).$
Next, we estimate $\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f\right\|_{L^{q}(\mu,{\mathbb{D}})}$. First, write
[…]
Again, denote
[…]
Therefore, using (3.3) with $E_{\alpha}$ replaced by $E_{4^{tk}}\backslash E_{4^{t(k+1)}}$, and $\left\{Q_{4^{tk},i}\right\}_{i\geq 1}$ by $\left\{\widetilde{Q}_{4^{tk},i}\right\}_{i\geq 1}$, respectively, we derive that
RHS of (3.21)
$\lesssim\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\frac{\mu(\widetilde{Q}_{4^{tk},i})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\left(\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}{\mathcal{M}}^{{\mathcal{D}},\sigma}\left(|f|\sigma^{-1}\right)(z)\cdot\sigma^{\frac{1}{p}}(z)\sigma^{\frac{1}{p^{\prime}}}(z)dA(z)\right)^{q}$
$\lesssim\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\frac{\mu(\widetilde{Q}_{4^{tk},i})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\left(\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)\right)^{\frac{q}{p}}\cdot\left(\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\sigma(z)dA(z)\right)^{\frac{q}{p^{\prime}}}$
$=\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\frac{\mu(\widetilde{Q}_{4^{tk},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{tk},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\cdot\left(\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)\right)^{\frac{q}{p}}$
(3.22) $\leq\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)\right)^{\frac{q}{p}}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\left(\frac{\mu(\widetilde{Q}_{4^{tk},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{tk},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\right)^{\frac{p}{p-q}}\right)^{\frac{p-q}{p}}.$
Now for the first double summation in (3.3), we have
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)$ $\lesssim\int_{{\mathbb{D}}}\left|{\mathcal{M}}^{{\mathcal{D}},\sigma}(|f|\sigma^{-1})(z)\right|^{p}\sigma(z)dA(z)$
$\lesssim\int_{{\mathbb{D}}}\left|f(z)\sigma^{-1}(z)\right|^{p}\sigma(z)dA(z)$
(3.23) $\lesssim\int_{{\mathbb{D}}}|f(z)|^{p}\omega(z)dA(z),$
where in the second estimate above, we have used the boundedness of ${\mathcal{M}}^{{\mathcal{D}},\sigma}$ acting on $L^{p}(\sigma,{\mathbb{D}})$ for any $1<p\leq+\infty$.
While for the second double summation in (3.3), we observe that
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\left(\frac{\mu(\widetilde{Q}_{4^{tk},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{tk},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{tk},i}|^{tq}}\right)^{\frac{p}{p-q}}$ $\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{tk},i}^{\textrm{up}}|\left(\frac{\mu(\widetilde{Q}_{4^{tk},i})\sigma^{\frac{q}{p^{\prime}}}(\widetilde{Q}_{4^{tk},i}^{\textrm{up}})}{|\widetilde{Q}_{4^{tk},i}|^{tq}|\widetilde{Q}_{4^{tk},i}|^{\frac{p-q}{p}}}\right)^{\frac{p}{p-q}}$
(3.24) $\lesssim\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textrm{up}}}\phi^{\frac{p}{p-q}}(z)dA(z)\leq\int_{{\mathbb{D}}}\phi^{\frac{p}{p-q}}(z)dA(z).$
Plugging (3.3) and (3.3) back to (3.3), we derive that
$\int_{{\mathbb{D}}}|{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)|^{q}\mu(z)dA(z)\lesssim\left\|f\right\|_{L^{p}(\omega,{\mathbb{D}})}^{q}\left\|\phi\right\|_{L^{\frac{p}{p-q}}({\mathbb{D}})},$
which gives $(2)\Longrightarrow(1)$.
$(1)\Longrightarrow(2)$.   Let $\{a_{I}\}_{I\in{\mathcal{D}}}$ be any sequence of positive numbers, and
$F(z):=\sum_{I\in{\mathcal{D}}}a_{I}\sigma(z)\mbox{1\hskip-4.25pt{1}}_{Q_{I}^{\textrm{up}}}(z).$
First, we note that
(3.25) $\int_{\mathbb{D}}|F(z)|^{p}w(z)dA(z)=\sum_{I\in{\mathcal{D}}}a_{I}^{p}\int_{Q_{I}^{\textrm{up}}}\sigma^{p}(z)w(z)dA(z)=\sum_{I\in{\mathcal{D}}}a_{I}^{p}\sigma(Q_{I}^{\textrm{up}})\lesssim\sum_{I\in{\mathcal{D}}}a_{I}^{p}\sigma(Q_{I}).$
On the other hand, for any $z\in Q_{I}^{\textrm{up}}$, one has
${\mathcal{M}}_{t}^{{\mathcal{D}}}F(z)\geq\frac{1}{|Q_{I}|^{t}}\int_{Q_{I}}|F(z)|dA(z)\geq\frac{1}{|Q_{I}|^{t}}\int_{Q^{\textrm{up}}_{I}}|F(z)|dA(z)=\frac{a_{I}\sigma(Q_{I}^{\textrm{up}})}{|Q_{j}|^{t}}\gtrsim\frac{a_{I}\sigma(Q_{I})}{|Q_{j}|^{t}},$
where in the last estimate above, we have used the assumption that $\sigma$ is dyadic doubling. Therefore,
$\int_{{\mathbb{D}}}\left|{\mathcal{M}}_{t}^{{\mathcal{D}}}F(z)\right|^{q}\mu(z)dA(z)$ $\gtrsim\sum_{I\in{\mathcal{D}}}\int_{Q_{I}^{\textrm{up}}}\left(\frac{a_{I}\sigma(Q_{I})}{|Q_{j}|^{t}}\right)^{q}\mu(z)dA(z)$
$\gtrsim\sum_{I\in{\mathcal{D}}}\frac{a_{I}^{q}\sigma^{q}(Q_{I})}{|Q_{I}|^{tq}}\mu(Q_{I}^{\textrm{up}})$
$\gtrsim\sum_{I\in{\mathcal{D}}}\frac{a_{I}^{q}\sigma^{q}(Q_{I})}{|Q_{I}|^{tq}}\mu(Q_{I})$
$=\sum_{I\in{\mathcal{D}}}\frac{\mu(Q_{I})\sigma^{\frac{q}{p^{\prime}}}(Q_{I})}{|Q_{I}|^{tq}}\cdot\left(a_{I}^{p}\sigma(Q_{I})\right)^{\frac{q}{p}}$
$=\sum_{I\in{\mathcal{D}}}\left(\beta_{I}|Q_{I}|^{1-\frac{q}{p}}\right)\cdot\left(a_{I}^{p}\sigma(Q_{I})\right)^{\frac{q}{p}},$
where in the third estimate above, we use the fact that $\mu\in{\bf B}_{\infty}({\mathcal{D}})$. Since ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}(\omega,{\mathbb{D}})\to L^{q}(\mu,{\mathbb{D}})$ is bounded, using the above estimate together with (3.25), we derive that
(3.26) $\sum_{I\in{\mathcal{D}}}\left(\beta_{I}|Q_{I}|^{1-\frac{q}{p}}\right)\cdot\left(a_{I}^{p}\sigma(Q_{I})\right)^{\frac{q}{p}}\lesssim\left(\sum_{I\in{\mathcal{D}}}a_{I}^{p}\sigma(Q_{I})\right)^{\frac{q}{p}},$
where the implicit constant in the above estimate is independent of the choice of $\{a_{I}\}_{I\in{\mathcal{D}}}$. Since $p>q$ and the choice of $\{a_{I}\}_{I\in\mathcal{D}}$ is arbitrary, duality yields that
$1\gtrsim\sum_{I\in{\mathcal{D}}}\left(\beta_{I}|Q_{I}|^{1-\frac{q}{p}}\right)^{\frac{p}{p-q}}\simeq\sum_{I\in{\mathcal{D}}}|Q_{I}^{\textrm{up}}|\beta_{I}^{\frac{p}{p-q}}=\left\|\phi\right\|_{L^{\frac{p}{p-q}}({\mathbb{D}})}^{\frac{p}{p-q}}.$
The proof is complete. ∎
@@ 4. Critical line estimates for the hypersingular Bergman projection
which can be regarded as a singular integral counterpart of the hypersingular maximal operator $M_{t}^{{\mathcal{D}}}$. We first recall that the off-critical line $L^{p}$ theory was studied in [14, Theorem 3]. In particular, they showed that for $K_{2t}$ is bounded from $L^{p}({\mathbb{D}})$ to $L^{q}({\mathbb{D}})$ if and only 1/q1/p>2t21/q-1/p>2t-2 for $1\leq p,q\leq\infty$, which is exactly the off-critical line regime for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$. Therefore, it is natural to consider the behavior of $K_{2t}$ on the critical line
@@
where in the last estimate, we used the standard integral estimate [37, Theorem 1.12]. Therefore, for any $\alpha>0$, as in Lemma 3.4,
@@
On the other hand, by [37, Theorem 1.12], one has
@@
First, recall that for any $z,w\in{\mathbb{D}}$, there exists a Carleson tent $Q_{J}$, containing both $z$ and $w$ such that $|Q_{J}|\simeq\left|1-z\overline{w}\right|^{2}$ (see, e.g., [31, Section 3]). Then using the $1/3$-trick, one can find some $I\in{\mathcal{D}}$ or $\widetilde{{\mathcal{D}}}$ (see, Remark 1.4), such that
@@
where ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ is the hypersingular sparse operator associated to a dyadic system $\mathcal{D}$} on $\mathbb{T}$, defined as [^4] We call ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ a hypersingular sparse operator since, on the critical line $\{(\tfrac{1}{p},\tfrac{1}{q}):\ \tfrac{1}{q}-\tfrac{1}{p}=2t-2\}$, it fails to satisfy strong $(p,q)$ bounds. Indeed, any strong-type estimate for ${\mathcal{A}}_{{\mathcal{S}},{\mathcal{D}}}^{t}$ at a point on this line would imply the corresponding bound for $K_{2t}$, which is known to be false by [14, Theorem 3].
@@ 5.1. Graded family
    Indeed, $K_{{\mathcal{G}}}^{\prime}$ only controls the ratios $\mathfrak{G}_{j}/\mathfrak{G}_{j+1}$ for sufficiently large $j$ and, in particular, imposes no restriction on the initial scales (for instance, on $\mathfrak{G}_{0}/\mathfrak{G}_{1}$). Consequently, one loses uniform control on the gaps between the first few layers. Using the same idea as in Observation 1.9, one can construct a sequence of sparse families with the same sparseness and the same value of $K_{{\mathcal{G}}}^{\prime}$, but for which the associated hypersingular sparse operators still exhibit the “blow-up” phenomenon described in Example 1.9.
@@
Theorem 5.7 strengthens our earlier results on the boundedness behavior of $K_{2t}$ (see, Section 4). Moreover, the proof of Step II in Theorem 5.7 yields an alternative proof of [14, Theorem 3]. Finally, we note that the endpoint weak-type estimate (5.3) fails in general (see, Proposition 4.2).
@@ 6. Some open problems
where ${\mathbb{B}}_{n}$ is the unit ball in $\mathbb{C}^{n}$ and $dV$ is the normalized volume measure on ${\mathbb{B}}_{n}$. In [34], Zhao and Zhou characterized the strong-type $L^{p}_{\alpha}({\mathbb{B}}_{n})\to L^{q}_{\beta}({\mathbb{B}}_{n})$ bounds for these operators under various assumptions on the parameters $a,b,c,\alpha$, and $\beta$. Here, for $1\leq p<\infty$ and $1<\alpha<\infty$, the space $L^{p}_{\alpha}({\mathbb{B}}_{n})$ is defined with respect to the measure $dv_{\alpha}(z)=c_{\alpha}(1-|z|^{2})^{\alpha}\,dV(z)$, where $c_{\alpha}$ is chosen so that $v_{\alpha}({\mathbb{B}}_{n})=1$. A natural question is whether the method developed in the present paper can be applied to obtain critical-line estimates for Forelli–Rudin type operators in the hypersingular regime (see, e.g., [34, Theorem 1.1]).